OpenStudy (anonymous):

The planning committee voted to increase the customer service budget by 10% from last year. If the budget last year was 56,000, what is the budget for the new year?

OpenStudy (bahrom7893):

The budget increased by 10%, the original one was 100%, so the new one is 110% (100+10 = 110) So let: 56,000 - 100% x - 110% Set up a proportion for x: x = (56,000 * 110%) / 100% = 560 * 110 = 61600

OpenStudy (bahrom7893):

wish my hw was like this right now.. =)

OpenStudy (anonymous):

hHAHAHA im hatting mines right now

OpenStudy (anonymous):

your in middle school?

OpenStudy (bahrom7893):

no im a freshie in college

OpenStudy (anonymous):

Lucky!! im hatting

OpenStudy (bahrom7893):

wait till u get to college....

OpenStudy (anonymous):

im i cant wait cause my teacher say its the funnest experience of your life

OpenStudy (bahrom7893):

well for me senior year in hs was the best year in my life!

OpenStudy (anonymous):

but if your not good at math your not going anywhere i herad is it true? and really ?

OpenStudy (bahrom7893):

it is, im a math major, got 5s on both AP calc ab and bc, and barely passed calc 3 in college, MTH4410 is killin me now

OpenStudy (anonymous):

So basically im not going to college or even passing if that stupid math think isnt working for me?

OpenStudy (bahrom7893):

not really, there's plenty of math involved, but thats not the only thing

OpenStudy (mathteacher1729):

Learning personal finance is CRITICAL in "real life". Beyond that, I'd say you should learn basic statistics because that will help keep you informed and learn how to spot lies and misdirection in the news. As for careers -- knowing math (specifically stats) opens up many many doors. Good computer programming skills can do the same, but they're built on a knowledge of math and analytical thinking. Long story short: Personal Finance = goal #1

OpenStudy (bahrom7893):

good point

OpenStudy (mathteacher1729):

Right now, the problems you're posting are very good practice for situations you'll run into in real life. Say you take a 2% cut in pay and you know that last year you made $36,897 (after taxes). How much do you make this year? Knowing how to solve this sort of thing is VERY important! Consider a happier problem. Say you want to buy a new car, and it costs $23,887. There is a 3% sales tax, and a 0.25% fee for the dealership. You know you can spend half the money you have saved in your on the down payment, which is 25% of the price of the car, before taxes. Right. See how complex it gets? Most adults CANNOT do this kind of math, and it hurts them, financially, VERY badly. The salesman can tell you ANYTHING they like, and if you can't check the math, you have no way of knowing if they are lying or not. So personal finance, learning percentages and stuff -- VERY VERY important! Get practice now, and you'll be thankful later. :)

OpenStudy (bahrom7893):

You know what guys? I know math is really hard. But stick to it! I barely passed AP Calculus AB (only passed one test in my junior year in hs), studied really hard over the summer took AP Calc BC next year, got 100 average, got 5s on both AB and BC, and skipped two semesters of calculus in college

OpenStudy (bahrom7893):

I know it sounds amazing, but it's not! If somebody told me I would get 5s in ab and bc back in junior year, i would've said ur lying

OpenStudy (bahrom7893):

and the stuff that im doing now is fun! Like in economics they will teach you the rule of 72, how 72/% = number of years it will take for ur money to double

OpenStudy (bahrom7893):

But if u actually do the math, the rule of 72 is sooo off!